# Task Diversity Shortens the In-Context Learning Plateau

### v2

@@ Abstract
Machine Learning, ICML
Figure 1: We train a transformer from scratch on in-context learning tasks. Single-task ICL: Training loss ( ) and test error/accuracy ( ) when each task is trained individually. The Parity task cannot be learned even after $1000$k training steps. Multi-task ICL: Training loss ( ) and test error/accuracy ( ) when all five tasks are trained simultaneously. Green lines mark the plateau escape points. Surprisingly, multi-task ICL training significantly shortens these plateaus, making training easier.
@@ 1.1 Related works
In-context learning (ICL) abilities of pretrained Large Language Models have gained significant attention since first investigated by GPT-3 (Brown et al., 2020). Following this, a large body of empirical studies have explored ICL in Large Language Models (Min et al., 2022a, b; Liu et al., 2021; Nie et al., 2022; Rubin et al., 2022; Wei et al., 2023). Given the complexity of real-world data, researchers have explored ICL in more stylized and simplified setups. Garg et al. (2022) formalized an approach to studying transformer’s performance to learn simple function class in context. Building on this work, several works have investigated the ability of models to learn stylized function classes in various settings, including boolean functions (Bhattamishra et al., 2024), regular language (Akyürek et al., 2024), as well as task mixtures (Tripuraneni et al., 2023), and the ability of Mamba (Park et al., 2024; Grazzi et al., 2024; Li et al., 2024). For more comprehensive overview, please refer to the survey by Dong et al. (2024).
@@
Multi-task training (Caruana, 1997; Baxter, 2000) is a widely used approach for model pretraining. In particular, researchers have identified that task diversity is crucial for pre-trained models to outperform in downstream tasks, alongside a large body of research across various domains: supervised learning (Tripuraneni et al., 2020, 2021; Du et al., 2021; Maurer et al., 2016; Crawshaw, 2020; Ruder, 2017), reinforcement learning (Zhang et al., 2023a; Jin et al., 2020; Hu et al., 2021; Yang et al., 2021; Collins et al., 2021; Lu et al., 2022; Cheng et al., 2022), and natural language processing (Zhang et al., 2023a; Zhao et al., 2023; Zhang et al., 2023b; Hu et al., 2020; Song et al., 2020; Zhou et al., 2019; Gunasekar et al., 2023; Sharma et al., 2023). These works mostly focused on statistical benefits, whereas our finding emphasize on the optimization benefits.
@@ 2 Experimental setup
Test loss. To evaluate the model’s performance on ICL tasks, we measure the error of the model’s prediction on the last ($n$-th) output of the prompt. For continuous ICL tasks, the test error is $(M_{\theta}(P_{n})-f(x_{n}))^{2}$. For boolean ICL tasks, the test accuracy is measured by $\mathbf{1}[\mathrm{sign}(M_{\theta}(P_{n}))=f(x_{n})]$, where $\mathbf{1}$ is the indicator function.
@@ 3 Task diversity shortens ICL plateaus
Table 1: Task diversity shortens the ICL plateau. We train a transformer with various combinations of $6$ different tasks ($d=10$). For each run, we report two metrics: the time to escape the plateau and the time to complete the learning of the task (written in parentheses). The rows correspond to the number of tasks trained together, and each entry in the table corresponds to the average time across the training runs that include the given task. For example, the entry at (Number of tasks = 3, Sparse Parity(2)) shows the average of $\binom{6}{3}$ results. We find that multi-task training shortens the ICL plateau. Precise details are provided in Appendix A.
@@
However, not all ICL tasks mutually reduce the duration of plateaus. For instance, we consider Regbench task (Akyürek et al., 2024), which is generated by a random automata. Our experiments in Appendix B.2 show that combining Regbench with the ICL tasks we consider does not reduce the plateau of Regbench, but does reduce the plateau of the other ICL tasks.
@@
Table 2: Muti-task ICL with uneven probabilities. We conducted experiments on multi-task ICL with uneven mixture probabilities, where $5$ ICL tasks are mixed with the ratio $\left(\frac{1}{2},\frac{1}{8},\frac{1}{8},\frac{1}{8},\frac{1}{8}\right)$. Under these uneven mixtures, we consistently observed a shortened plateau.
@@
1 $\infty$ $\infty$ $2.8$k ($4.2$k) $10.2$k ($11.5$k) $2.3$k ($3.3$k)
5 $2.4$k ($2.9$k) $3.0$k ($6.8$k) $2.0$k ($2.5$k) $2.4$k ($4.1$k) $2.0$k ($2.5$k)
Figure 3: Mamba (First row) and Hyena (Second row). Red lines and Blue lines respectively represent the loss dynamics of single-task training and multi-task training. (Left): Quadratic Regression vs. Quadratic Regression$+$Linear Regression. (Middle): Sparse Parity(2) vs. Sparse Parity(2)$+$Linear Regression. (Right): Sparse Parity(3) vs. Sparse Parity(3)$+$Quadratic Regression.
@@
In the ICL setups we consider, we observe that (i) there is only one plateau, and (ii) learning is very rapid once this plateau is escaped from. This implies that the time at which training is completed, defined as the moment when the model reaches near-perfect training accuracy, is very close to the time of escaping from the (first) plateau. The results of Table 1 confirm this.
@@ 4.1 Plateau is task-wise no-context learning
i.e., $g_{\mathcal{F}}^{\star}$ is the context-independent function that minimizes the test error. For boolean ICL tasks, $g_{\mathcal{F}}^{\star}$ is analogously defined with the argmax of the test accuracy. In many cases, the optimal no-context function has a simple closed-form expression. For instance, if =$\mathcal{F}=\{f\,|\,f(x)=w^{\intercal}x\}$ and $\mathcal{D}_{\mathcal{F}}$ is given by $w\sim\mathcal{N}(\mu,I_{d})$, then $g_{\mathcal{F}}^{\star}(x)=\mu^{\intercal}x$. As another example, if =$\mathcal{F}=\{f\,|\,f(x)=x^{\intercal}Wx\}$ and $\mathcal{D}_{\mathcal{F}}$ is given by $W\sim\mathcal{N}(U,I_{d\times d})$, then $g_{\mathcal{F}}^{\star}(x)=x^{\intercal}Ux$. The ICL plateau corresponds to task-wise no-context learning, which we describe in the following. When we train a model $M_{\theta}$ for ICL with function classes $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$, the model’s output during its plateau corresponds to
@@
In other words, the model identifies the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ and then applies the optimal no-context function corresponding to $\mathcal{F}_{m}$. The in-context demonstrations $(x_{1},f(x_{1}),\dots,x_{n-1},f(x_{n-1}))$ are used to determine the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ but not to determine the specific $f\in\mathcal{F}_{m}$. This claim can be verified by measuring the error between $M_{\theta}(P_{n})$ and $g^{\star}_{\mathcal{F}}$, which we plot in Figure 5. Revisiting the Sparse Parity(2) task, the accuracy during plateau is indeed attributed to $0.55=\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}[\mathbf{1}(g_{\mathcal{F}}^{\star}(x)=f(x))]$. Appendix E provides further details on no-context learning regime.
@@ 4.2.1 Checkpoint experiment
The findings, summarized in Figure 6, indicate that the checkpoint model transferred from task A quickly learns task B with a shortened plateau. This implies that (1) task A and task B share a common structure and that (2) the plateau arises from the difficulty of learning the common structure. We conduct an analogous experiment on natural language ICL tasks and observe similar results, as shown in Figure 8 of the appendix. Further details are presented in Appendix C.
@@ 4.2.2 Generality of our hypothesis
Figure 7: (Left) Language ICL task. Previous work Fu et al. (2024) identified the difficulty of learning the WordSelection$(4)$ task. We found that mixing it with WordLength or WordSelection$(2)$ reduces the plateau. Refer to Appendix B.1 for further details. (Right) Feature learning setup. For the toy model described in Section 4.2.2, multi-task feature learning significantly shortens the loss plateau. Refer to Appendix D for further details.
@@ 4.2.3 The common structure is not just an induction head
Retrieval. Sample $1024$ 5-tuples of one key and four values: $\{(k_{i},v_{i}^{1},v_{i}^{2},v_{i}^{3},v_{i}^{4})\}_{i=1}^{1024}$. The $k_{i}$s and $v_{i}^{j}$s are independently sampled from $\mathcal{D}_{\mathcal{K}}$ and $\mathcal{D}_{\mathcal{V}}$, respectively. To generate prompts, we uniformly sample $(n-1)$ 5-tuples without replacement and uniformly choose one $v_{i}$ per 5-tuple, resulting in $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{n-1}$. Next, we sample $p\sim\mathrm{Unif\,}(\{1,\dots,n-1\})$ and set $\mathbf{q}=\mathbf{k}_{p}$. Finally, given $P_{n}=(\mathbf{k}_{1},\mathbf{v}_{1},\dots,\mathbf{k}_{n-1},\mathbf{v}_{n-1},\mathbf{q})$, the task is to predict $\mathbf{v}_{p}$. We consider two Retrieval tasks: Gaussian Retrieval with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathcal{N}(0,I_{d}),\mathcal{N}(0,1))$ and Boolean Retrieval with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathrm{Unif\,}\{\pm 1\}^{d},\mathrm{Unif\,}\{\pm 1\})$.
Note that the induction head is precisely the mechanism for solving this Retrieval task. We conduct a checkpoint experiment for the Retrieval tasks, allowing the models to learn the induction head. To train continuous ICL tasks, we use a checkpoint model pre-trained with Gaussian Retrieval. To train Boolean ICL tasks, we use a checkpoint model pre-trained with Boolean Retrieval. We find that the checkpoint does not significantly shorten the ICL tasks’ plateaus. This result demonstrates that the common structure shared by other ICL tasks is not just an induction head. (It is unclear whether an induction head is useful for our ICL tasks at all.) For further details, please refer to Appendix C.2.
@@ 5 In-context learning tasks
  • Linear Regression. =$\mathcal{F}=\{f\,|\,f(x)=w^{\intercal}x\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
  • Quadratic Regression. =$\mathcal{F}=\{f\,|\,f(x)=x^{\intercal}Wx\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $W\in\mathbb{R}^{d\times d}$ is independently sampled from $\frac{1}{\sqrt{d}}\mathcal{N}(\mu,1)$.
  • Sparse Linear Regression. =$\mathcal{F}=\{f\,|\,f(x)=w_{\text{sparse}}^{\intercal}x\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$. To sample $w_{\text{sparse}}$, we uniformly choose $k=3$ coordinates and retain the corresponding coordinates of $w$.
  • LeakyReLU Regression. =$\mathcal{F}=\{f\,|\,f(x)=\mathrm{LeakyReLU\,}(w^{\intercal}x)\}$ with negative slope $\alpha=0.5$. $\mathcal{D}_{\mathcal{F}}$: Each elements of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
@@
  • Sparse Parity$(k)$. =$\mathcal{F}=\{f\,|\,f(x)=\prod_{i\in A}x[i]\}$. $\mathcal{D}_{\mathcal{F}}$: $A\subseteq\{1,\dots,d\}$ is a uniformly sampled subset of size $k$.
  • Parity. =$\mathcal{F}=\{f\,|\,f(x)=\prod_{i\in A}x[i]\}$. $\mathcal{D}_{\mathcal{F}}$: $A\subseteq\{1,\dots,d\}$ is a uniformly sampled subset, regardless of the size.

### v3

@@ 1 Introduction
Figure 1: We train a transformer from scratch on in-context learning tasks. Single-task ICL: Training loss ( ) and test error/accuracy ( ) when each task is trained individually. The Parity task cannot be learned even after 10001000k training steps. Multi-task ICL: Training loss ( ) and test error/accuracy ( ) when all five tasks are trained simultaneously. Green lines mark the plateau escape points. Surprisingly, multi-task ICL training significantly shortens these plateaus, making training easier.
@@ 1.1 Related works
In-context learning (ICL) abilities of pretrained Large Language Models have gained significant attention since first investigated by GPT-3 (Brown et al., 2020). Following this, a large body of empirical studies have explored ICL in Large Language Models (Min et al., 2022a; b; Liu et al., 2021; Nie et al., 2022; Rubin et al., 2022; Wei et al., 2023). Given the complexity of real-world data, researchers have explored ICL in more stylized and simplified setups. Garg et al. (2022) formalized an approach to studying transformer’s performance to learn simple function class in context. Building on this work, several works have investigated the ability of models to learn stylized function classes in various settings, including boolean functions (Bhattamishra et al., 2024), regular language (Akyürek et al., 2024), as well as task mixtures (Tripuraneni et al., 2023), and the ability of Mamba (Park et al., 2024; Grazzi et al., 2024; Li et al., 2024). For more comprehensive overview, please refer to the survey by Dong et al. (2024).
@@
Multi-task training (Caruana, 1997; Baxter, 2000) is a widely used approach for model pretraining. In particular, researchers have identified that task diversity is crucial for pre-trained models to outperform in downstream tasks, alongside a large body of research across various domains: supervised learning (Tripuraneni et al., 2020; 2021; Du et al., 2021; Maurer et al., 2016; Crawshaw, 2020; Ruder, 2017), reinforcement learning (Zhang et al., 2023a; Jin et al., 2020; Hu et al., 2021; Yang et al., 2021; Collins et al., 2021; Lu et al., 2022; Cheng et al., 2022), and natural language processing (Zhang et al., 2023a; Zhao et al., 2023; Zhang et al., 2023b; Hu et al., 2020; Song et al., 2020; Zhou et al., 2019; Gunasekar et al., 2023; Sharma et al., 2023). These works mostly focused on statistical benefits, whereas our finding emphasize on the optimization benefits.
@@ 2 Experimental setup
Test loss. To evaluate the performance on ICL tasks, we measure the model’s prediction error on the last ($n$-th) output of the prompt. For continuous tasks, the test error is $(M_{\theta}(P_{n})-f(x_{n}))^{2}$. For boolean tasks, the test accuracy is $\mathbf{1}[\mathrm{sign}(M_{\theta}(P_{n}))=f(x_{n})]$, where $\mathbf{1}$ is the indicator function.
@@ 3 Task diversity shortens ICL plateaus
Table 1: Task diversity shortens the ICL plateau. We train a transformer with various combinations of 66 different tasks ($d=10$). For each run, we report two metrics: the time to escape the plateau and the time to complete the learning of the task (written in parentheses). The rows correspond to the number of tasks trained together, and each entry in the table corresponds to the average time across the training runs that include the given task. For example, the entry at (Number of tasks = 3, Sparse Parity(2)) shows the average of $\binom{6}{3}$ results. We find that multi-task training shortens the ICL plateau. To further robustly support our claims, we repeated each experiment three times with varying random seeds and reported the mean values. Details are provided in Appendix C.
@@
Figure 3: Mamba (First row) and Hyena (Second row). Red lines and Blue lines respectively represent the loss dynamics of single-task training and multi-task training. (Left): Quadratic Regression vs. Quadratic Regression++Linear Regression. (Middle): Sparse Parity(2) vs. Sparse Parity(2)++Linear Regression. (Right): Sparse Parity(3) vs. Sparse Parity(3)++Quadratic Regression.
@@
However, not all ICL tasks mutually reduce the duration of plateaus. For instance, we consider Regbench task (Akyürek et al., 2024), which is generated by a random automata. Our experiments in Appendix D.2 show that combining Regbench with the ICL tasks we consider does not reduce the plateau of Regbench, but does reduce the plateau of the other ICL tasks.
@@
Table 2: Muti-task ICL with uneven probabilities. We conducted experiments on multi-task ICL with uneven mixture probabilities, where 55 ICL tasks are mixed with the ratio $\left(\frac{1}{2},\frac{1}{8},\frac{1}{8},\frac{1}{8},\frac{1}{8}\right)$. Under these uneven mixtures, we consistently observed a shortened plateau.
@@
1 $\infty$ $\infty$ 2.82.8k (4.24.2k) 10.210.2k (11.511.5k) 2.32.3k (3.33.3k)
5 2.42.4k (2.92.9k) 3.03.0k (6.86.8k) 2.02.0k (2.52.5k) 2.42.4k (4.14.1k) 2.02.0k (2.52.5k)
@@
In the ICL setups we consider, we observe that (i) there is only one plateau, and (ii) learning is very rapid once this plateau is escaped from, as illustrated in Appendix 8. This implies that the time at which training is completed, defined as the moment when the model reaches near-perfect training accuracy, is very close to the time of escaping from the (first) plateau. The results of Table 1 confirm this.
@@ 4.1 Plateau is task-wise no-context learning
i.e., $g_{\mathcal{F}}^{\star}$ is the context-independent function that minimizes the test error. For boolean ICL tasks, $g_{\mathcal{F}}^{\star}$ is analogously defined with the argmax of the test accuracy. In many cases, the optimal no-context function has a simple closed-form expression. For instance, if $\mathcal{F}=\{f\,|\,f(x)=w^{\intercal}x\}$ and $\mathcal{D}_{\mathcal{F}}$ is given by $w\sim\mathcal{N}(\mu,I_{d})$, then $g_{\mathcal{F}}^{\star}(x)=\mu^{\intercal}x$. As another example, if $\mathcal{F}=\{f\,|\,f(x)=x^{\intercal}Wx\}$ and $\mathcal{D}_{\mathcal{F}}$ is given by $W\sim\mathcal{N}(U,I_{d\times d})$, then $g_{\mathcal{F}}^{\star}(x)=x^{\intercal}Ux$. The ICL plateau corresponds to task-wise no-context learning, which we describe in the following. When we train a model $M_{\theta}$ for ICL with function classes $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$, the model’s output during its plateau corresponds to
@@
In other words, the model identifies the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ and then applies the optimal no-context function corresponding to $\mathcal{F}_{m}$. The in-context demonstrations $(x_{1},f(x_{1}),\dots,x_{n-1},f(x_{n-1}))$ are used to determine the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ but not to determine the specific $f\in\mathcal{F}_{m}$. This claim can be verified by measuring the error between $M_{\theta}(P_{n})$ and $g^{\star}_{\mathcal{F}}$, which we plot in Figure 5. Revisiting the Sparse Parity(2) task, the accuracy during plateau is indeed attributed to $0.55=\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}[\mathbf{1}(g_{\mathcal{F}}^{\star}(x)=f(x))]$. Appendix G provides further details on no-context learning regime.
@@ 4.2.1 Checkpoint experiment
The findings, summarized in Figure 6, indicate that the checkpoint model transferred from task A quickly learns task B with a shortened plateau. This implies that (1) task A and task B share a common structure and that (2) the plateau arises from the difficulty of learning the common structure. We conduct an analogous experiment on natural language ICL tasks and observe similar results, as shown in Figure 15 of the appendix. Further details are presented in Appendix E.
@@ 4.2.2 Generality of our hypothesis
Figure 7: (Left) Language ICL task. Previous work Fu et al. (2024) identified the difficulty of learning the WordSelection(4)(4) task. We found that mixing it with WordLength or WordSelection(2)(2) reduces the plateau. Refer to Appendix D.1 for further details. (Right) Feature learning setup. For the toy model described in Section 4.2.2, multi-task feature learning significantly shortens the loss plateau. Refer to Appendix F for further details.
@@ 4.2.3 The common structure is not just an induction head
Retrieval. Sample $1024$ 5-tuples of one key and four values: $\{(k_{i},v_{i}^{1},v_{i}^{2},v_{i}^{3},v_{i}^{4})\}_{i=1}^{1024}$. The $k_{i}$s and $v_{i}^{j}$s are independently sampled from $\mathcal{D}_{\mathcal{K}}$ and $\mathcal{D}_{\mathcal{V}}$, respectively. To generate prompts, we uniformly sample (n1)(n-1) 5-tuples without replacement and uniformly choose one $v_{i}$ per 5-tuple, resulting in $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{n-1}$. Next, we sample $p\sim\mathrm{Unif\,}(\{1,\dots,n-1\})$ and set $\mathbf{q}=\mathbf{k}_{p}$. Finally, given $P_{n}=(\mathbf{k}_{1},\mathbf{v}_{1},\dots,\mathbf{k}_{n-1},\mathbf{v}_{n-1},\mathbf{q})$, the task is to predict $\mathbf{v}_{p}$. We consider two Retrieval tasks: Gaussian Retrieval with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathcal{N}(0,I_{d}),\mathcal{N}(0,1))$ and Boolean Retrieval with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathrm{Unif\,}\{\pm 1\}^{d},\mathrm{Unif\,}\{\pm 1\})$.
Note that the induction head is precisely the mechanism for solving this Retrieval task. We conduct a checkpoint experiment for the Retrieval tasks, allowing the models to learn the induction head. To train continuous ICL tasks, we use a checkpoint model pre-trained with Gaussian Retrieval. To train Boolean ICL tasks, we use a checkpoint model pre-trained with Boolean Retrieval. We find that the checkpoint does not significantly shorten the ICL tasks’ plateaus. This result demonstrates that the common structure shared by other ICL tasks is not just an induction head. (It is unclear whether an induction head is useful for our ICL tasks at all.) For further details, please refer to Appendix E.2.
@@ 5 In-context learning tasks
  • Linear Regression. $\mathcal{F}=\{f\,|\,f(x)=w^{\intercal}x\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
  • Quadratic Regression. $\mathcal{F}=\{f\,|\,f(x)=x^{\intercal}Wx\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $W\in\mathbb{R}^{d\times d}$ is independently sampled from $\frac{1}{\sqrt{d}}\mathcal{N}(\mu,1)$.
  • Sparse Linear Regression. $\mathcal{F}=\{f\,|\,f(x)=w_{\text{sparse}}^{\intercal}x\}$. $\mathcal{D}_{\mathcal{F}}$: Each element of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$. To sample $w_{\text{sparse}}$, we uniformly choose $k=3$ coordinates and retain the corresponding coordinates of $w$.
  • LeakyReLU Regression. $\mathcal{F}=\{f\,|\,f(x)=\mathrm{LeakyReLU\,}(w^{\intercal}x)\}$ with negative slope $\alpha=0.5$. $\mathcal{D}_{\mathcal{F}}$: Each elements of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
@@
  • Sparse Parity(k)(k). $\mathcal{F}=\{f\,|\,f(x)=\prod_{i\in A}x[i]\}$. $\mathcal{D}_{\mathcal{F}}$: $A\subseteq\{1,\dots,d\}$ is a uniformly sampled subset of size $k$.
  • Parity. $\mathcal{F}=\{f\,|\,f(x)=\prod_{i\in A}x[i]\}$. $\mathcal{D}_{\mathcal{F}}$: $A\subseteq\{1,\dots,d\}$ is a uniformly sampled subset, regardless of the size.